# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v4

@@ 2.4 Boundary conditions and connection formulae
with ${\cal F}_{\rm inst}(a,x)$ defined in (21). Collecting all the different pieces together, we have at our disposal four different bases of solutions, so that the general solution can be written as
$G_{1}(z)$ $=$ $\sum_{\alpha}c_{\alpha}g_{\alpha}(x)G^{0}_{\alpha}({\textstyle\frac{x}{z}})+\ldots=\sum_{\alpha,\alpha^{\prime}}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}({\textstyle\frac{x}{z}})+\ldots$ (38)
$=$ $\sum_{\alpha}c_{\alpha}G^{1}_{\alpha}(z)+\ldots=\sum_{\alpha,\alpha^{\prime}}c_{\alpha}F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}(z)+\ldots$
The first line corresponds to the PM approximation where $x$ is taken small keeping $y$ finite. The second equality in this line is obtained using the connection formulae (26). Similarly, the second line corresponds to the PN approximation where $x$ is taken small keeping $z$ finite and the right hand side follows from (35). Finally dots stand for terms of order $x^{k+1}$ in the corresponding limits. Expanding at the leading order the two right hand sides in (38) and using (20), one finds the asymptotics
@@ 3.1 Heun gravity correspondence
$R(z)$ $=$ $e^{-{x\over 2z}}(1-z)^{1-{m_{1}+m_{2}\over 2}}({\textstyle\frac{x}{z}})^{1+m_{3}}G_{1}(z)\qquad,\quad z={2M\over r}\qquad,\qquad x=4iM\omega$ (44)
@@
of the fundamental solutions $R_{\alpha}(y)$. Since we are mainly interested in the physics at large distances $z\ll 1$, it is convenient to work with the PM representation and write the two independent solutions of the Teukolsky equation as
@@ 4 Waveform and energy flux at 30PN
In this Section we use perturbation theory to compute the probe limit of the wave form and the energy flux produced by a circular EMRI binary system at order 30PN, i.e. $v^{60}$, with $v$ the tangential velocity of the light particle. The results will be compared against the 22PN results available in the literature Fujita:2012cm and the numerical computations using the package Teukolsky of the Black Hole Perturbation Toolkit (BHTK) BHPToolkit based on the MST method Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk and on a numerical integrations of the differential equation. We stress the fact that by 30PN we mean that the polynomials $P_{0}$, $\widehat{P}_{0}$ and the exponent $a$, defining the solution, are truncated at order 30PN, but the hypergeometric functions and the $c^{\rm in}_{\alpha}(a)$ coefficients specifying the solution are kept exact, so an infinite number of terms beyond 30PN are taken into account. In particular, the infinite tower of tail contributions (log divergences, $\pi$’s and transcendental numbers) are resummed into exponentials and Gamma functions. The observables we consider are the $(\ell,m)$ harmonic modes of the incoming solution $\mathfrak{R}_{\text{in}}$ and the energy flux radiated to infinity by the system.
@@ 4.2 The luminosity
${}_{s}Y_{\ell m}(\theta,\phi)$ $=$ $e^{im\phi}\sin^{2\ell}\left(\frac{\theta}{2}\right)\sqrt{\frac{(2\ell{+}1)(\ell{-}m)!(\ell{+}m)!}{4\pi(\ell{-}s)!(\ell{+}s)!}}$ (84)
$\sum_{r=0}^{\ell{-}s}({-})^{\ell{+}m{+}r{-}s}{{\ell-s}\choose{r}}{{\ell+s}\choose{r+s-m}}\cot({\textstyle\frac{\theta}{2}})^{2r{+}s{-}m}$
@@ 4.3 Numerical tests
The results are compared in Table 1 against those obtained with the MST method and numerical integration using the BHTK BHPToolkit code. In all the subsequent tables, the displayed digits are those that are insensitive to the contribution of terms of order higher than $v^{60}$ in $P_{0}(y)$ and $\widehat{P}_{0}(y)$. The agreement is up to 10 and 14 digits even for small values of $r$, where the PM approximation is less reliable.
@@
Figure 2: Luminosity vs velocity. We display the $v$-dependence of the ratio between $\eta_{\ell m}$ and the Newton luminosity (87) for the three dominating modes. a: (2,2), b: (3,3), c: (2,1).
@@
  • RinRupdRdt.nb: with this file it is possible to compute the $\eta_{\ell m}$ modes importing the data files P30.m, Pt30.m, gg30.m, gk30.m and a30.m, containing the polynomials $P_{0}(y)$, $\widehat{P}_{0}(y)$, the quantities $\gamma$, $\gamma_{\kappa}$ and the $a$ cycle. The results thus obtained will show a slower convergence with respect to the data of Tables 1, 2-3 and 4 which were obtained keeping higher orders than 30PN for $c_{ij},\widehat{c}_{ij}$. The files with these data are quite large and can be obtained by writing to any of the authors.
@@ 5 Summary and conclusions
In this paper, using black hole perturbation theory, we have derived exact formulae for the PMk and PNk expansions (at order $k$ in the soft-limit $\omega M\ll 1$) of the gravitational wave form emitted by a particle moving on an arbitrary trajectory in the Schwarzschild geometry at leading order in the probe limit. The results rely on a novel hypergeometric representation of the solutions of the Heun (or confluent Heun) equation (see Appendix A for the discussion of the non-confluence case). We build four equivalent basis of solutions, $\widetilde{G}^{1}_{\alpha}$, $G^{1}_{\alpha}$, $G^{0}_{\alpha}$ and $\widetilde{G}^{0}_{\alpha}$ that describe with good accuracy different regions of the space time, and are connected to each other by connection matrices. The final picture is summarized in the following scheme
@@
We remark that our results provide a hypergeometric representations of solutions of the Heun equations of general type, so they apply to any physical problem governed by equations of Heun type. In particular, they can be easily adapted to the study of collisions of spinning BHs (Kerr geometry) or orbital motions in the backgrounds of topological stars, D-branes and cosmological perturbations Bianchi:2021xpr ; Bianchi:2021mft ; DiRusso:2024hmd ; Bianchi:2024vmi ; Bianchi:2024mlq ; Bianchi:2024rod . It would also be very interesting to apply our results to the generation of templates of waveforms for GWs data analysis [^7] We thank A.Nagar and M.Panzeri for useful discussions on this issue.. We hope to also come back to this issue in a future publication.

### v5

@@ 2.4 Boundary conditions and connection formulae
with ${\cal F}_{\rm inst}(a,x)$ defined in (22). Collecting all the different pieces together, we have at our disposal four different basis of solutions $\{G^{0}_{\alpha},\widetilde{G}^{0}_{\alpha},G^{1}_{\alpha},\widetilde{G}^{1}_{\alpha}\}$ related to each other by the connection matrices $B_{\alpha\alpha^{\prime}}$, $F_{\alpha\alpha^{\prime}}$. More precisely, the general solution can be written as
$G_{1}(z)$ $=$ $\sum_{\alpha}c_{\alpha}g_{\alpha}(x)G^{0}_{\alpha}({\textstyle\frac{x}{z}})=\sum_{\alpha,\alpha^{\prime}}c_{\alpha}g_{\alpha}(x)B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}({\textstyle\frac{x}{z}})$ (39)
$=$ $\sum_{\alpha}c_{\alpha}G^{1}_{\alpha}(z)=\sum_{\alpha,\alpha^{\prime}}c_{\alpha}F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}(z)$
The first line corresponds to the PM approximation where $x$ is taken small keeping $y$ finite. The second equality in this line is obtained using the connection formulae (27). Similarly, the second line corresponds to the PN approximation where $x$ is taken small keeping $z$ finite and the right hand side follows from (36). We stress that these connection formulae are exact. As in the case of the hypergeometric connection formulae , they can be written as ratios of gamma functions with the caveat that the arguments of these functions depend on $a(u)$. The whole non-triviality of the confluent Heun equation is codified in a single function $a(u)$. Expanding at the leading order the two right hand sides in (39) and using (21), one finds the asymptotics
@@ 3.1 Heun gravity correspondence
$R(z)$ $=$ $e^{-{x\over 2z}}(1-z)^{2{-}{x\over 2}}({\textstyle\frac{x}{z}})^{{-}1{-}{x\over 2}}G_{1}(z)\qquad,\quad z={2M\over r}\qquad,\qquad x=4iM\omega$ (45)
@@
of the two independent solutions of the Teukolsky equation, let us say in the PM representation, that according to (45) are given by
@@ 4 Wave form and energy flux at 30PN
In this Section we use perturbation theort to compute the probe limit of the wave form and the energy flux produced by a circular EMRI binary system at order 30PN, i.e. $v^{60}$, with $v$ the tangential velocity of the light particle. The results will be compared against the 22PN results available in the literature Fujita:2012cm and the numerical computations using the package Teukolsky of the Black Hole Perturbation Toolkit (BHTK) BHPToolkit based on the MST method Tagoshi:1993dm ; Poisson:1994yf ; Shibata:1994jx ; Tagoshi:1994sm ; Tanaka:1996lfd ; Mano:1996vt ; Mino:1997bx ; Suzuki:1998vy ; Fujita:2010xj ; Fujita:2011zk and on a numerical integrations of the differential equation. We stress the fact that by 30PN, we mean that the polynomials coefficients $P_{0}$, $\widehat{P}_{0}$ and the exponent $a$, defining the solution are truncated at order 30PN, but the hypergeometric functions and the $c^{\rm in}_{\alpha}(a)$ coefficients specifying the solution are kept exact, so an infinite number of terms beyond 30PN are taken into account. In particular, the infinite tower of tail contributions (log divergences, $\pi$’s and transcendental numbers) are resummed into exponentials and Gamma functions. The observables we consider are the $(\ell,m)$ harmonic modes of the incoming solution $\mathfrak{R}_{\text{in}}$ and the energy flux radiated to infinity by the system.
@@ 4.2 The luminosity
${}_{s}Y_{\ell m}(\theta,\phi)=e^{im\phi}\sin^{2\ell}\left(\frac{\theta}{2}\right)\sqrt{\frac{(2\ell{+}1)(\ell{-}m)!(\ell{+}m)!}{4\pi(\ell{-}s)!(\ell{+}s)!}}\sum_{r=0}^{\ell{-}s}({-})^{\ell{+}m{-}r{-}s}(^{\ell{-}s}_{s})(^{\ell+s}_{r{+}s{-}m})\cot({\textstyle\frac{\theta}{2}})^{2r{+}s{-}m}$ (85)
@@ 4.3 Numerical tests
The results are compared in Table 1 against those obtained by the MST method and numerical integration using the BHTK BHPToolkit . In all the subsequent tables, the displayed digits are those ones that are insensible to the contribution of terms of order higher than $v^{60}$ in $P_{0}(y)$ and $\widehat{P}_{0}(y)$. The agreement is up to 10 and 14 digits even for small values of $r$, where the PM approximation is less reliable.
@@
Figure 2: Luminosity vs velocity. We display the $v$-dependence of the ratio $\eta_{\ell m}/\eta^{0}_{22}$ for the three dominating modes. a: (2,2), b: (3,3), c: (2,1).
@@
  • RinRupdRdt.nb With this file it is possible to compute the $\eta_{\ell m}$ modes importing the data files P30.m, Pt30.m, gg30.m, gk30.m and a30.m, containing the polynomials $P_{0}(y)$, $\widehat{P}_{0}(y)$, the quantities $\gamma$, $\gamma_{\kappa}$ and the $a$ cycle. The results thus obtained will show a slower convergence with respect to the data of Tables 1, 2-3 and 4 which were obtained keeping higher orders than 30PN for $c_{ij},\widehat{c}_{ij}$. The files with these data are quite large and can be obtained by writing to any of the authors.
@@ 5 Summary and conclusions
In this paper, using black hole perturbation theory, we derive exact formulae for the PMk and PNk expansions (at order $k$ in the soft-limit $\omega M\ll 1$) of the gravitational wave form emitted by a particle moving on an arbitrary trajectory in Schwarzschild geometry at leading order in the probe limit. The results rely on a novel hypergeometric representation of the solutions of the Heun (or confluent Heun) equation (see Appendix A for the discussion of the non-confluence case). We build four equivalent basis of solutions, $\widetilde{G}^{1}_{\alpha}$, $G^{1}_{\alpha}$, $G^{0}_{\alpha}$ and $\widetilde{G}^{0}_{\alpha}$ that describe with good accuracy different regions of the space time, and are connected to each other by Heun connection matrices. The final picture is summarized in the following scheme
@@
We remark that our results provide a hypergeometric representations of solutions of Heun equations of general type, so they apply to any physical problem governed by equations of Heun type. In particular, they can be easily adapted to the study of collisions of spinning BHs (Kerr geometry) or orbital motions in the backgrounds of topological stars, D-branes and cosmological perturbations Bianchi:2021xpr ; Bianchi:2021mft ; DiRusso:2024hmd ; Bianchi:2024vmi ; Bianchi:2024mlq ; Bianchi:2024rod . It would also be very interesting to apply our results to the generation of templates of waveforms for GWs data analysis [^7] We thank A.Nagar and M.Panzeri for useful discussions on this issue.. We hope to also come back to this issue in a future publication.